OpenStudy (superhelp101):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 29 ? 155 159 (i think this one) 163 167

OpenStudy (aum):

\(\large a_n = a_1 + (n-1)d\) \(\large a_6 = 4 + (6-1)d\). Find d. Then plug it into: \(\large a_{32} = 4 + (32-1)d\)

OpenStudy (superhelp101):

ok.. give meh a sec to find it :))

OpenStudy (superhelp101):

yes...I got 159

OpenStudy (aum):

correct.

OpenStudy (superhelp101):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -35 ? -89 -86 -83 -80

OpenStudy (superhelp101):

let me solve it and let ya know ;) just to check

OpenStudy (superhelp101):

I keep getting -48/16 :/

OpenStudy (superhelp101):

oh never mind that's for d.

OpenStudy (superhelp101):

I got -89 :)

OpenStudy (aum):

correct.

OpenStudy (superhelp101):

yay! :)

OpenStudy (superhelp101):

What is the sum of a 34-term arithmetic sequence where the first term is 14 and the last term is 146? 2,550 2,720 2,890 2,982 I am going to find it and let ya know in a sec. Just to make sure..